OpenStudy (fibonaccichick666):

PROVE: Let f be riemann integrable on [a,b] Let \(\alpha , \beta , \gamma \) be arbitrary numbers in [a,b] (not necessarily ordered in any way) Prove that \[\int^\gamma_\alpha f=\int^\beta_\alpha f+ \int^\gamma_\beta\]

OpenStudy (fibonaccichick666):

just a hint here, I don't see how it is different from the thm 5.2.2 from last time

OpenStudy (zzr0ck3r):

consider the case with \(\beta>\gamma\)

OpenStudy (fibonaccichick666):

right, but wouldn't there be 4 potential options?

OpenStudy (zzr0ck3r):

yeah I guess so, you would have really 6.....

OpenStudy (fibonaccichick666):

beta greater than gamma is just that thm from the last problem right?

OpenStudy (zzr0ck3r):

post that link again. I thought it was nested.

OpenStudy (fibonaccichick666):

nested?

OpenStudy (fibonaccichick666):

http://www.jirka.org/ra/realanal.pdf page157

OpenStudy (zzr0ck3r):

it says b<c in that theorem

OpenStudy (fibonaccichick666):

oh i guess i meant the \(\gamma<\beta\) option

OpenStudy (zzr0ck3r):

yes that would be taken care of by that theorem

OpenStudy (fibonaccichick666):

said it backwards shoot

OpenStudy (zzr0ck3r):

so you have some rule about \(\int_a^bf=-\int_b^af\)?

OpenStudy (fibonaccichick666):

yea, we have that

OpenStudy (zzr0ck3r):

I guess you will have six cases then? \(a\le b\le c\\ a\le c\le b\\ b\le a \le c\\ b \le c \le a\\ c\le a \le b\\ c \le b \le a\)

OpenStudy (fibonaccichick666):

so we need cases \(\gamma<\alpha ; \beta<\alpha ; \beta>\alpha; \beta>\gamma\)

OpenStudy (fibonaccichick666):

we already have \(\beta<\gamma\), but do I have to do all of them? like 6! solutions?

OpenStudy (fibonaccichick666):

there must be some general way

OpenStudy (zzr0ck3r):

how do you have \(\beta \le \gamma\)?

OpenStudy (fibonaccichick666):

we can discredit equals because that equals zero...

OpenStudy (zzr0ck3r):

how do you have \(\beta < \gamma\)?

OpenStudy (fibonaccichick666):

that thm from before

OpenStudy (zzr0ck3r):

5.2.2?

OpenStudy (zzr0ck3r):

it does not say that.

OpenStudy (fibonaccichick666):

why not? Didn't we prove that?

OpenStudy (zzr0ck3r):

lol im confused. we did not prove theorem 5.2.2 it is in your book. you said that theorem 5.2.2 says that we must have beta < gamma, and then you say we proved it. I think I am missing something.

OpenStudy (zzr0ck3r):

brb, getting off of bus and will get on pc when I get home. Should not take long.

OpenStudy (fibonaccichick666):

ok thanks, I'm so confused on what I need to prove here

OpenStudy (fibonaccichick666):

I guess I also don't see how this works when we have \[\int^\gamma_\alpha f=\int^\beta_\alpha f+ \int_\gamma^\beta\]

ganeshie8 (ganeshie8):

Maybe prove a lemma first and use it instead of working all the cases : \[\int \limits_{\alpha}^{\beta}f = -\int\limits_{\beta}^{\alpha} f\]

OpenStudy (fibonaccichick666):

right but how is that the same?

OpenStudy (fibonaccichick666):

when we subtract that? ... ohhhh ok I'm a nitwit, ok, so I just need to prove that now. Let me try that

OpenStudy (fibonaccichick666):

so conceptually I get it, we are subtracting the excess because we have modified our values, but can I get a hint on the technical proof?

ganeshie8 (ganeshie8):

proving lemma should be straightforward right ?

ganeshie8 (ganeshie8):

you're asking how to use it to argue that the integral splitting does not depend on the order ?

OpenStudy (fibonaccichick666):

hmm ok... I'll try let me hit my head off the wall for a few minutes and I'll be back

ganeshie8 (ganeshie8):

i don't get what you're asking

OpenStudy (fibonaccichick666):

what you consider straightforward, is not to me

ganeshie8 (ganeshie8):

okay its straightforward in calc1 atleast idk how it changes in real analysis

OpenStudy (fibonaccichick666):

so I'm gonna try and see what I'm missing

ganeshie8 (ganeshie8):

\[\int_a^b f = F(b)-F(a) = -\left[F(a) - F(b)\right] = -\int_b^a f \]

OpenStudy (fibonaccichick666):

oh that lemma, yea that has already been proved, I thought you meant the subtracion of the integral to cut off the excess

ganeshie8 (ganeshie8):

yeah making that connection is tricky, one sec

OpenStudy (fibonaccichick666):

but first, do we have 6 pieces or do we have more?

OpenStudy (fibonaccichick666):

so like if we use \(\gamma<\alpha\) do we then have 4 sub-cases where \(\beta <\alpha ; \beta<\gamma; \beta>\gamma; \beta >\alpha\)

OpenStudy (fibonaccichick666):

which means we have 24 cases??? That seems like alot

ganeshie8 (ganeshie8):

there should be a clever way to argue instead of listing out all the different cases

ganeshie8 (ganeshie8):

i see only 6 cases though

OpenStudy (fibonaccichick666):

hmm

ganeshie8 (ganeshie8):

the ones listed by zzr

OpenStudy (fibonaccichick666):

yea, but isn't there multiple subcases if you bound it?

OpenStudy (fibonaccichick666):

ie the limits are gamma and alpha, so you have two possibilities there

ganeshie8 (ganeshie8):

those 6 cases exhaust all the subcases

OpenStudy (fibonaccichick666):

either gamma is bigger or alpha is bigger, then there are 4 subcases

OpenStudy (fibonaccichick666):

beta is bigger/less than alpha, beta is bigger/less than gamma

OpenStudy (fibonaccichick666):

all four for both options on bounds

ganeshie8 (ganeshie8):

list ur cases based on the "middle value" Case1 : when alpha is in between beta and gamma, you get two subcases \[\beta \le \alpha \le \gamma \] \[\gamma \le \alpha\le \beta\]

OpenStudy (fibonaccichick666):

but the end values are not set either

ganeshie8 (ganeshie8):

yes you get two mose cases with two subcases each ?

ganeshie8 (ganeshie8):

*more

OpenStudy (fibonaccichick666):

how am I getting 10...

ganeshie8 (ganeshie8):

im not sure, actually im bit lost...

ganeshie8 (ganeshie8):

think of it like this : you have 3 points on real line 

ganeshie8 (ganeshie8):

you can permute them in 3! = 6 ways

OpenStudy (fibonaccichick666):

... I abhor analysis. Thank god next week is the last week of it....

OpenStudy (fibonaccichick666):

I am thinking 2 options 4 suboptions

ganeshie8 (ganeshie8):

somehow your extra cases must be redundant

OpenStudy (fibonaccichick666):

hmm ok,

OpenStudy (fibonaccichick666):

ah I see the redundant ones now

ganeshie8 (ganeshie8):

Oh good, im still trying to prune the redundant ones

OpenStudy (fibonaccichick666):

but still, 6 cases seems like a lot, shouldn't there be another way?

ganeshie8 (ganeshie8):

So you work all 6 cases by subtracting the excess where necesssary ?

ganeshie8 (ganeshie8):

yeah im fully sure there will be another nice way but im not getting any other ideas as of nw

OpenStudy (fibonaccichick666):

yea... but I also don't see how to technically and rigorously prove that

ganeshie8 (ganeshie8):

@zzr0ck3r is still off

OpenStudy (fibonaccichick666):

Conceptually I fully understand it

ganeshie8 (ganeshie8):

same here idk real analysis terminology

OpenStudy (fibonaccichick666):

it's the technical..... ughh